UT 20150508 LUCI2 AO Nighttime Commissioning

Observer LUCI-Team: JHeidt, APramskiy, DThompson
Observer AO: DMiller (LBTO)
Telescope Operator: SAllenson
LBTO Support: TSargent

Plan for the night

- Continue AO-com depending on observing conditions

- May do first AFC-tests on sky

Summary

  • lost 2/3 of night due to weather

  • Progress on AFC, close to go on sky in ima-mode

  • LS-acq. in AO-mode works

Details

All times are UT
Files are luci2.20150506.NNNN.fits

02:10 Sunset, mostly cloudy, some scattered showers around, T = 1deg, Hum = 85%,
strong and gusty winds (15-20m/s). Should clear up by 11pm or so, need to wait a bit

Field alignment: FM4/M1 -10100 steps, FM4/M2 +1800 steps

06:00 Still closed, cloudy, humidity and wind close to closing limit (85-90% and 15-17m/s)

08:00 Wind and humidity are below limit, we give it a try, Steve is going to open

0089 0090 test exposures to check for condensation. OK so far

08:34 L2_Acqui_BS9177_PA80.xml
modified to use cam=1
  1. (sky) 0093 (source) 0094 (slit)
    Acq offsets are calculated for N3.75 camera...because that is what LUCI thinks is in!
    Calculated by hand assuming 0.015"/pixel and submitted w/TelescopeService
  2. post-offset image...why is the mask back in turnout?
    star @ 923.98 1018.05
    Insert mask in FPU
  3. BrG image. Star is in slit
  4. K image. Actually looks decently centered

08:57 L2_Acqui_BS9177_PA0.xml
modified to use cam=1
  1. (sky) 0099 (source) 0100 (slit)
    Offset 1.900 -4.147" calculated by hand
    moving mask to turnout for post-offset image
  2. post-offset image, star @ 877.11 1023.51
  3. with mask in FPU, BrG
  4. K image
  5. first spectrum - nothing obvious
  6. second spectrum

  1. 0110 diagnosis why we don't see anything...G210 used...v.high resolution
  2. star still in slit!

Moving to running Observer panel from luci.luci (SW mis-match on obs2?)

09:43 L2_Acqui_BS9177_PA0.xml - Re-submit script
  1. (sky) 0113 (source) 0114 (slit)
    Script behaving oddly, integration time seems wrong for this step

09:56 L2_Acqui_BS9177_PA0.xml - Re-submit script after copying back to luci.luci
  1. (sky) 0116 (source) 0117 (slit)
    Manual offset -2.109 -4.197"
  2. (source frame, mask got moved out!) @ 925.42 1019.00
    Move mask back in FPU
  3. - 0121 (source+slit) slit @ 921.79 (savemode was NORMAL)
    offset -0.054 in X
  4. -0124 (source+slit) (savemode was normal)

  1. 0127 G040 @ 2.17um I don't see anything
  2. 0129 G040 @ 1.65um still nothing (no sky lines either)
  3. 0132 Check with N3.75 camera...wavelength calibration is off, H spec @ right edge of FoV

10:33 BS9177_ghost_K.xml
modified to use cam=1
  1. xx 01xx 100 x 3.0s LIR/NORMAL !!! (no data taken)
    forced abort (pause + turned off instrument and readout)

10:41 BS9177_ghost_K.xml
changed savemode to INTEGRATED
  1. 0134 100 x 3.0 LIR/INT
    G040+J @ 1.25 ??? prior abort did not cleanly end?

10:56 re-load BS9177_ghost_K.xml
Re-load from disk, reset, refresh...instr. config worked this time but
the status panel at right still shows G040@1.25+K though it is definitely mirror
JH suggests to restart the scheduler
  1. 0136 100 x 3.0s LIR/INT
    star near lower left edge in 0135, but offset is -10 -10

11:18 Do ZP_N30_rest.xml
  1. Nope...95% humidity, closing

Suggestions:
On queue panel:
reset and refresh buttons should make a visible change in the queue.
"skip" label should say "skip to here"
"warning" should be "Mask in Turnout"
should be able to abort whatever is happening (wanted to do this before RunAO)

Click/unclick instrument tab not cleanly handled?

-- %USERSIG{JochenHeidt - 2015-05-01}%
Topic revision: r7 - 08 May 2015, JochenHeidt
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback